### v1

# On the extraction of fission mode properties from fragment mass distributions

Abstract
Background: Fission modes are typically characterised by fragment mass and total kinetic energy centroids, around which a distribution of these variables is observed. These distributions are usually fitted with Gaussian functions.
Purpose: To investigate how the properties of these “Gaussian fission modes” compare with underlying “theoretical fission modes” defined from the potential energy surface of the fissioning nuclei.
Methods: A simple approach, inspired by the scission point model, is introduced to investigate the impact of anharmonicity of the potential along the scission line on Gaussian mode properties. This approach is also used to evaluated an “effective potential” from the yields.
Results: Several Gaussian functions are usually required to fit yields from non-harmonic potentials associated with a unique theoretical mode. Similarly, “effective fission modes”, defined from the wells of the effective potentials, are sometime very different to the Gaussian modes. For instance, the S1 and S2 Brosa modes contribute to the same effective potential well at scission.
Conclusions: Gaussian fits are useful to identify the role of shell effects in fission. However, comparisons with theoretical potential energy surfaces are better carried with effective potentials extracted from the yields, assuming that a broad range of excitation energies is available.
I Introduction
Nuclear fission allows for a sensitive dynamic probe of nuclei and the behaviour of nucleons within them. Even the earliest measurements of the mass and charge yields of fission fragments [1, 2] identified a distribution of masses, with some nuclides tending to fission asymmetrically, and others symmetrically. Labelling these as fission modes swiftly followed [3]. The symmetric and asymmetric modes exhibit total kinetic energy (TKE) distributions characteristic to the charge asymmetry as well as the elongation of the pre-scission nucleus [4]. The origin of these modes is nowadays interpreted in terms of shell effects in the fissioning nucleus [5, 6, 7, 8] as well as in the fragments [9, 10, 11, 12, 13, 8]. In particular, Brosa et al [14] tied the fission modes to particular pre-scission shapes. Although they approximated the distribution of observables from each mode as being a normal or Gaussian distribution, they also asserted that they “do not consider the Gaussian to be more than an approximate representation of the yield from one fission channel” and note that experimental evidence for the existence of distinct modes requires detailed measurement of the change in yields with small changes in excitation energy and neutron number, to reveal the bifurcations between these channels.
In the 21st century, continual improvements in accelerator and detector technology have allowed for more and more measurements of fission fragment mass and TKE, revealing a variety of fission modes. Many publications do not explicitly fit particular modes to their data [15, 16, 17, 18, 19, 20, 21, 22, 23, 24], however those that do almost invariably fit the fragment mass (or charge) distributions using a collection of Gaussians [25, 26, 27, 28, 29, 30, 31, 32, 33]. Assuming that a symmetric mode leads to a Gaussian distribution centred around symmetric mass split, and that each asymmetric mode is associated with a pair of Gaussians (centred around the most likely heavy and light fragments for this mode), leads to the interpretation of multiple modes based on the quality of these fits. However, the underlying assumption of the Gaussian mode is unquestioned, despite the fact that theoretical models like the scission point model [9] or the time-dependent generator coordinate method (TDGCM) [34, 35] usually do not predict yields with Gaussian shapes. Nevertheless, Gaussian fits of yield distributions are useful to identify the role of shell effects through the stability of Gaussian centroids with respect to changes in excitation energy and fissioning nucleus. However, it is not clear how these “Gaussian modes” translate into features of the underlying potential energy surface (PES). Indeed, the yield distributions could in principle be used to constrain the shape of the PES near to the scission line [36, 37, 38, 39, 40, 41, 42].
The purpose of this work is to evaluate the relationship between fission mode properties extracted from Gaussian fits of fragment mass or charge distributions, and the properties of the potential energy of the fissioning system near scission, such as fission valleys in PES. In particular, we introduce a simple expression to extract an “effective potential energy” that can be used to compare experimental yields with theoretical PES properties near scission, provided that a broad range of excitation energies is considered.
Expressions used to parametrise fragment mass asymmetries and yields, as well as the various definitions of fission modes used in this work, are provided in Sec. II.1. Section III discusses the impact of the potentials on the yields and their Gaussian fits properties. Effective potentials are extracted from Gaussian fits of experimental data, and their dependence with excitation energies are analysed in Sec. IV. Conclusions are drawn in Sec. V.
II Model
II.1 Parametrisation of fragments mass asymmetry
For simplicity, we assume that there is a single possible TKE for a given mass split. This allows us to parameterise yields with one coordinate associated with mass repartition between fragments. We define the coordinate $x$ to compare the mass ratio of the fragments in a way that allows easy fitting and comparison, and is unitless:
$x=4\frac{A_{1}-A_{T}/2}{A_{T}},$ (1)
where $A_{1}$ is the mass number of one fragment and $A_{T}$ is the mass number of the compound nucleus. By definition, $x$ ranges from $-2$ to $2$, however the region of interest is generally contained within $[-1,1]$. Note that, under the unchanged charge density assumption, $x$ can also be used to compare charge ratio of the fragments.
II.2 Parametrisation of yield distribution
Inspired by the scission point model [9] (see also [43, 44, 45, 46, 47, 48, 49, 50]), we write the yield as
$Y(x)=e^{-\frac{V(x)}{T}}.$ (2)
In the scission point model, i.e., under the assumption of quasiequilibrium among collective degrees of freedom near the scission point, $V(x)$ is the potential energy of a shape associated with $x$ immediately before scission, while $T$ is the temperature of the collective system. Equation (2) will also be used as a way to parametrise the yield, in which case $V(x)$ will be treated as an “effective potential” used as a proxy for the actual potential along the scission line. In particular, the effective potentials will be used to interpret the Gaussian fits of the yield in terms of potential shapes. The “temperature” $T$ is used to account for a dependence with the excitation energy $E^{*}$ of the form $T=\sqrt{E^{*}/a}$, with $a=10$ MeV-1. The yield normalisation only leads to a constant shift of $V(x)$ and is irrelevant. Other models which consider random walks [51, 52] or diffusion paths [53, 54, 55, 40, 56] or even coherent wavepacket evolution [57, 58, 59, 35, 60] along a potential energy surface are not exactly equivalent to this simple picture. Nevertheless, in the end, the topography of the potential energy is always responsible for the resulting shape of the yield distribution.
II.3 Definitions of fission modes
The definition of a fission mode (or fission channel) is not necessarily the same from a theoretical or experimental perspective. Experimentally, fission modes are associated with structures in the fragment mass, charge, and TKE distributions, usually fitted with Gaussian functions. In particular, different modes correspond to different Gaussian centroids. Theoretically, fission modes are associated with valleys in the PES, or, in the context of the scission point model, minima in the scission line. In this case, different modes correspond to different local minima separated by potential energy ridges or barriers.
As the potential energy is in principle not an observable, the connection between both characterisations of fission modes is not straightforward. To facilitate this connection, we propose an alternative definition of fission modes based on the effective potential $V(x)$. The latter can be evaluated directly from the yield distribution using Eq. (2). The minima of $V(x)$ (which is a symmetric function) then indicate the degree of asymmetry of the fission modes. A symmetric mode is associated with a local minimum at $x=0$, while an asymmetric mode is associated with two local minima [^1] In principle, the barrier between the minima should be at least of the order of the zero point energy. at $x=\pm x_{0}\neq 0$.
In agreement with a quantum treatment that predicts Gaussian ground-states of harmonic oscillators, a symmetric Gaussian yield would lead to a harmonic effective potential $V(x)\propto x^{2}$. Similarly, at low enough excitation energy, one could expect asymmetric wells in $V(x)$ to produce a Gaussian yield centred around ±$\pm x_{0}$ due to the approximate quadratic shape of the potential near its minima at ±$\pm x_{0}$. As will be discussed in details in this work, however, yields that require several Gaussians for a reasonable fit do not necessarily lead to one minimum per Gaussian in the effective potential, suggesting that the theoretical and experimental definitions of fission modes are not equivalent.
For clarity, let us summarise the three definitions of fission modes that are used in this work:
  • Gaussian modes are associated with the characteristics of the Gaussian functions used to fit the yield distribution;
  • theoretical modes correspond to valleys in the PES, or in the context of the scission point model, minima of potential energy along the scission line;
  • minima in the effective potential $V(x)$ [see Eq. (2)] are called effective modes.
In Sec. III, we analyse the yields induced by potentials within a simple approach based on the scission point model, that is then used to compare theoretical, Gaussian, and effective modes. In Sec. IV, Gaussian modes obtained from fits of experimental fission data are used to evaluate the effective potential, and then to compare Gaussian and effective modes.
III From potential to yield
In this section we investigate the shape of yield distributions induced by simple non-harmonic potentials $V(x)$. Here, $V(x)$ can be interpreted as describing the potential energy along the scission line, rather than being an effective potential. Following the scission point model, the yields are obtained from Eq. (2) assuming a quasiequilibrium among collective degrees of freedom near the scission point. This approximation neglects pre-scission dynamics and does not account for the effect of, e.g., large barriers that could prevent the system to explore a particular fission valley. Nevertheless, this simple model is sufficient to illustrate the difference between mode characterisations.
III.1 Symmetric mode
Figure 1: Quartic potential $V=x^{4}$ (black solid line) in MeV, and associated yield $Y(x)$ from Eq. (2) with $T=0.5$ MeV (red solid line). (a) Two- and (b) three-Gaussian fits (thick dashed lines). Thin dashed lines show each individual Gaussian component.
As mentioned in the previous section, a quadratic potential $V\propto x^{2}$ naturally leads to a yield with a symmetric peak (i.e., centred at $x=0$) and a Gaussian shape, in which case the Gaussian mode is a perfect representation of the theoretical mode. In a polynomial expansion of the potential, however, the quadratic term could also vanish, in which case one would expect a yield distribution deviating from a Gaussian shape. Consider, as an example, a quartic potential $V=x^{4}$ as represented in Fig. 1. The associated yield $Y(x)$ from Eq. (2) with $T=0.5$ MeV is clearly not Gaussian. Figs. 1 (a) and (b) also show two- and three-Gaussian fits to $Y(x)$. The quality of the fit is poor with two Gaussians, but gets much better with three Gaussians.
Figure 2: Centroids of the asymmetric Gaussian modes obtained from two- and three-Gaussian fits as a function of $T$ for a quartic potential.
The centroids of the asymmetric Gaussian mode vary significantly with excitation energy, as shown in Fig. 2, despite the fact that the underlying potential remains the same. Based on the better agreement obtained with the three-Gaussian fit, one concludes that the system exhibits both a symmetric and an asymmetric Gaussian modes. However, the underlying physics is clear as there is only one symmetric theoretical mode in this system. The dependence of the centroid of the asymmetric Gaussian mode with excitation energy is also an indication that it is not associated with a theoretical mode.
Figure 3: Comparison between the quartic potential (solid line) and the potentials extracted from the two (blue dashed line) and three (red dashed line) Gaussian fits in Fig. 1 using Eq. (3).
The Gaussian fits of the yield
$Y_{G}(x)=\sum_{n=1}^{N_{G}}y_{n}e^{-\frac{(x-x_{n})^{2}}{2{\sigma_{n}}^{2}}}=Ce^{-\frac{V_{G}(x)}{T}},$
with $N_{G}$ the number of Gaussians, can also be used to evaluate the associated effective potential
$V_{G}(x)=-T\left[\ln\left(Y_{G}(x)\right)-\ln C\right],$ (3)
where $C$ is a constant. The resulting potentials are plotted in Fig. 3. Although the potential obtained from the two-Gaussian fit leads to a spurious asymmetric effective mode, the three-Gaussian fit correctly predicts a potential with a unique symmetric effective mode. This demonstrates that, for Gaussian fits to be compared with theoretical modes, they should be used to extract an associated effective potential, as in Eq. (3). In particular, the number of theoretical modes and their characteristics are different than those of Gaussian modes. Furthermore, in the framework of the scission point model, theoretical and effective modes have similar characteristics, providing the fit of the yield is of good quality. In the present case, this requires a minimum of three Gaussians (one symmetric and one asymmetric Gaussian modes). Note that further increasing the number of Gaussian modes would not change the conclusions as this would have a small impact on the effective potential.
Of course, one is not required to use Gaussians at all. Indeed, the potential can be extracted directly, e.g., using a polynomial expansion $V(x)=\sum_{n=1}^{N_{p}}a_{n}x^{2n}$, where $N_{p}$ is the degree of the polynomial, and fitting the yield with
$Y(x)=C\exp\left(-\frac{1}{T}\sum_{n=1}^{N_{p}}a_{n}x^{2n}\right)$ (4)
to extract the coefficients $a_{n}$.
III.2 Asymmetric mode
Figure 4: Two-well potential $V(x)=\frac{1}{4}-x^{2}+x^{4}$ (black solid line) in MeV, and yields $Y(x)=e^{-V(x)/T}$. Values of $T$ are in MeV.
Figure 4 shows an example of a two-well potential associated with one asymmetric theoretical mode with $x_{0}=1/\sqrt{2}$. Yields obtained at various excitation energies are also shown. As expected, the lower the excitation energy, the closer the peaks are to a Gaussian shape. At higher $T$, however, the increase of the yield at symmetry would require an additional symmetric Gaussian mode, despite the fact that the potential does not exhibit any symmetric theoretical mode.
Figure 5: Two-well potential $V=\frac{1}{4}-x^{2}+x^{4}$ (black solid line) and associated yield $Y(x)$ from Eq. (2) with $T=0.5$ MeV (red solid line). (a) Two- and (b) three-Gaussian fits (thick dashed lines). Thin dashed lines show each individual Gaussian component.
This is illustrated in Fig. 5(a) and (b) showing the yield at $T=0.5$ MeV together with two- and three-Gaussian fits, respectively. The two-Gaussian fits is of poor quality and in particular does not reproduce the peak position well. Adding a third Gaussian at symmetry considerably improves the fit. However, the symmetric Gaussian mode should not be misinterpreted as a symmetric theoretical mode.
Figure 6: Centroids of the asymmetric Gaussian modes obtained from two- and three-Gaussian fits as a function of $T$ for the two-well potential of Fig. 4 whose theoretical mode is indicated by the dotted line.
The excitation energy dependence of the position of the asymmetric Gaussian modes is shown in Fig. 6. Both two- and three-Gaussian fits exhibit significant variations and globally fail to predict the exact asymmetry with precision. Again, the excitation energy dependence of the Gaussian modes illustrates their difference with theoretical modes.
Figure 7: Comparison between the two-well potential (solid line) and the effective potentials extracted from the two (blue dashed line) and three (red dashed line) Gaussian fits in Fig. 5 using Eq. (3).
Figure 7 shows the effective potentials extracted from the Gaussian modes using Eq. (3). Here, both two- and three-Gaussian fits lead to a unique asymmetric effective mode. In particular, the symmetric Gaussian mode in the three-Gaussian fit does not translate into an additional symmetric effective mode. Nevertheless, the resulting effective potentials exhibit significant differences with the original two-well potential. As in the quartic case, the two-well potential can be exactly reproduced from a fit of the yield as given in Eq. (4).
IV From yield to effective potential
We now illustrate the method of extracting and analysing effective potentials from experimental yields. We consider three systems spanning a range of excitation energies: fusion-fission [61] and $\beta-$delayed fission [62] of 180Hg, as well as neutron induced fission of 238U and 232Th [25]. Rather than using experimental data as a starting point, we use the Gaussian fits of the fragment mass distributions as obtained by the authors. While 180Hg is well described by one asymmetric Gaussian mode, the actinides require two asymmetric and one symmetric Gaussian modes, also known as Standard 1 (S1), Standard 2 (S2), and symmetric superlong Brosa modes, respectively [14].
IV.1 180Hg
180Hg fission fragment mass distributions were reported in [62] for $\beta-$delayed fission of 180Tl experiments and in [61] for fusion-fission of 36Ar$+^{144}$m (see also [63] for similar reactions), spanning a broad range of excitation energies up to $E^{*}=65.5$ MeV. These yields were well reproduced by two Gaussian fits [61] that are reproduced in Fig. 8. The position of the Gaussian centroids remains relatively constant over the excitation energies [61], a fact that can be attributed to the influence of shell effects on yield distributions.
Figure 8: Two Gaussian fits (corresponding to one asymmetric Gaussian mode) of 180Hg experimental fission yields. The lowest excitation energy is from $\beta-$delayed fission [62] while the others are from 36Ar$+^{144}$m fusion-fission reactions [61].
Figure 9 shows that the effective potentials associated with these yields vary significantly with excitation energy. In particular, they take the shape of a two-well potential at low energy, but transition to a single potential well at higher energy.
Figure 9: 180Hg effective potentials from the yields of Fig. 8.
IV.2 238U neutron induced fission
Figure 10: Gaussian modes (dashed lines) and corresponding total yields (solid lines) obtained from fits of experimental fragment mass distributions resulting from 238U neutron induced fission at various neutron energies [25].
Gaussian modes obtained from 238U neutron-induced fission data [25] are shown in Fig. 10 for various neutron energy ranges. The S2 mode, which is the most asymmetric Gaussian mode, clearly dominates the asymmetric peaks. However, the increase of the yield at symmetry with increasing neutron energy indicates a stronger contribution of a symmetric Gaussian mode at higher excitation energy. The latter is very broad and span a range of $x$ well into the other modes.
Figure 11: 238U effective potentials from the total yields of Fig. 10. The average neutron energy was used to evaluate the excitation energies.
Figure 11 shows the effective potentials at excitation energies corresponding to the average neutron energy for each range. Each effective potential exhibits a clear two-well structure indicating an asymmetric effective mode. The asymmetry of this mode remains constant with excitation energy, as expected from an influence of shell effect. We also note a decrease of the effective potential barrier at symmetry with increasing excitation energy.
IV.3 232Th neutron induced fission
Figure 12: Gaussian modes (dashed lines) and corresponding total yields (solid lines) obtained from fits of experimental fragment mass distributions resulting from 232Th neutron induced fission at various neutron energies [25].
A similar analysis of Gaussian modes for neutron induced fission on 232Th at various neutron energies [25] is presented in Fig. 12 . The same S1, S2 and symmetric Gaussian modes are used in the fits. As in the case of 238U, the asymmetric peak is dominated by the S2 mode. At higher energy, however, the symmetric Gaussian mode becomes significant enough to induce a peak at symmetry. The effective potentials shown in Fig. 13 indeed indicate the appearance of a well at symmetry at high $E^{*}$.
Figure 13: 232Th effective potentials from the total yields of Fig. 12. The average neutron energy was used to evaluate the excitation energies.
V Discussion and conclusion
Despite the fact that anharmonic potentials may induce a change of Gaussian mode centroids with excitation energy, it is remarkable that these remain constant with excitation energy for all three nuclei studied in section IV. This illustrates the usefulness of Gaussian modes to identify the influence of shell effects on yield distributions. The role of shell effects on 180Hg fission fragment formation has been studied in several theoretical works [52, 64, 38, 65, 66, 67, 68, 12, 69]. In actinides, the S1 and S2 mode are centred around $Z\approx 52.5$ and $Z\approx 55$ protons in the heavy fragment, respectively [70]. A possible interpretation of the underlying shell effects is that they originate from octupole deformed shell gaps at $Z=52$ and $Z=56$ [11]. Indeed, near scission, the fragments acquire pear shapes induced by the competition between the short range attraction in the neck and the long range repulsion between them. The octupole deformed shell effects then facilitate these deformations. The $Z=50$ spherical shell effect, however, leads to a rapidly increasing energy with octupole deformation, thus acting against the formation of such fission fragments. Nevertheless, the latter induces a symmetric fission mode in neutron rich fermium isotopes (“supershort” Brosa mode [14]) where both fragments approach the doubly magic 132Sn nucleus.
Despite their different characteristics, the S1 and S2 asymmetric Gaussian modes produce a single asymmetric effective mode. The position of the latter remains stable with excitation energy, confirming its shell effect origin. This observation is in agreement with theoretical PES which, to our knowledge, do not exhibit different valleys for S1 and S2 Gaussian modes (see, e.g., [34]), indicating that they both contribute to the same “Standard” theoretical mode. Naturally, the shell effects responsible for S1 and S2 Gaussian modes are expected to induce a non-harmonic shape of the potential energy at scission. Note that the supershort Gaussian mode in neutron-rich fermium, attributed to $Z=50$ spherical shell effects, is associated with a different theoretical mode than the standard one, i.e., two valleys separated by a potential ridge are usually found in theoretical PES [71, 8].
The fusion-fission 180Hg data extend to relatively high excitation energies, for which a transition from an asymmetric to a symmetric effective mode is observed. The fact that the effective potential varies with excitation energy is not surprising as the shell effects are expected to weaken with increasing excitation energy, inducing the system to transition from a two-well to a one-well effective potential.
The 238U data show an increase of the symmetric Gaussian mode and a decrease of the barrier separating the asymmetric wells, with increasing excitation energy. This could also be interpreted in terms of reduction of shell effects. The case of 232Th, however, shows that an additional symmetric well in the effective potential appears at high excitation energy. A possible interpretation is that the system now has enough energy to overcome a barrier in the PES leading to a symmetric fission valley. The fact that a symmetric well is not observed in 238U data could be due to a higher saddle point preventing the system to enter the symmetric valley, which is often predicted by theory (see, e.g., [34]). This is an obvious limitation of using the effective potential to compare with theoretical predictions when only a limited range of excitation energies is available. Although, in principle, the comparison is valid within the framework of the scission point model, it neglects the pre-scission dynamics and non-trivial PES topography. The latter require more advanced approaches such as TDGCM [34, 35].
Summarising, Gaussian fits of fission fragment mass or charge distributions are useful to identify the role of shell effects in fragment formation. However, these “Gaussian fission modes” (as, e.g., the S1 and S2 modes) often contribute to the same potential well at scission (or fission valley in the PES), which are associated with “theoretical fission modes”. We proposed a simple way to convert fragment yields into an effective potential at scission, that can then be used to define “effective fission modes”. Assuming that a broad range of excitation energies is available, these effective modes provide an easy way to compare with theoretical modes predictions.